If f(x) = 2x + 3 and g(x) = x^2 + 1, find f(g(3))
Answer by amarjeeth123(574)

f(x) = 2x + 3
g(x) = x^2 + 1
g(3)=3^2+1=9+1=10
f(g(3))=f(10)=20+3=23.